The mean height of grade 9 students at a certain high school is 164 centimeters and the standard deviation is 10 centimeters. Assuming the heights are normally distributed, What percent of the heights is greater than 168 centimeters? 


1
Expert's answer
2021-02-28T07:21:46-0500

"P(X>168)=P(Z>\frac{168-164}{10})=P(Z>0.4)=0.3446=34.46\%."